# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

### v2

@@ 5.5 Quantitative Results
Main Results. Table 1 presents the performance comparison across all evaluated models. Our method achieves substantial improvements over baseline models on both datasets. On PubMedQA, we obtain 61.95% accuracy and 50.88% Macro F1, outperforming the best baseline (Gemma 7B) by 3.87% and 2.86%, respectively. On MedQA-USMLE, our approach reaches 27.92% accuracy and 20.41% Macro F1, surpassing the strongest baseline (MedAlpaca 7B) by 3.91% accuracy. Remarkably, our 1.1B parameter model achieves better overall performance than much larger models, including 7B and 13B variants, demonstrating the effectiveness of our reasoning refinement approach.
Ablation Studies. Table 2 analyzes the contribution of key components in our framework. The comparison between "No Refinement + RPRO" and "RPRO (with Refinement)" reveals that reasoning refinement provides substantial improvements, particularly for Macro F1 (38.11% → 50.88%), while maintaining similar accuracy levels. This suggests that refinement enhances the quality of reasoning beyond mere correctness. Comparing different training paradigms, RPRO with refinement outperforms both SFT and DPO variants, validating the effectiveness of group-wise preference optimization for multi-candidate reasoning scenarios.
@@
Figure 2: Performance comparison on PubMedQA and MedQA-USMLE under different $\beta$ values. Accuracy (solid blue line) and Macro F1 (dashed orange line) are reported.
In Fig. 2, we investigate the impact of different regularization coefficients $\beta$ on performance, revealing that moderate $\beta$ values can enhance Macro F1 while maintaining accuracy, thereby achieving a balance between factual correctness and reasoning comprehensiveness. Excessively low $\beta$ values cause the model to deviate from the reference distribution and lose stability, while excessively high $\beta$ values over-constrain the model and suppress reasoning diversity. These findings validate the critical importance of appropriate KL regularization for enhancing robustness and generalization capabilities in clinical reasoning tasks.
@@
Figure 3: Performance across different confidence thresholds on (a) PubMedQA and (b) MedQA-USMLE datasets. Accuracy (solid blue line) and Macro F1 (dashed purple line) are reported.
@@
Threshold and Convergency Analysis. Figure 3 examines the impact of the acceptance probability threshold $\tau$ on model performance. Lower thresholds (0.2, 0.4) result in excessive refinement, potentially degrading reasoning quality through over-correction. The optimal threshold of 0.6 balances the necessity for refinement with the preservation of reasoning. Higher thresholds (0.8) show marginal accuracy improvements but decreased Macro F1, suggesting that conservative refinement may miss opportunities for quality enhancement while maintaining correctness.
@@
Figure 4: Training loss curves on MedQA-USMLE (left) and PubMedQA (right) datasets. The plots show KL divergence, pairwise loss, ranking loss, and total loss across training steps.
@@
To evaluate the robustness of our training, we further analyze our RPRO training loss as shown in Fig. 4. The results demonstrate that RPRO exhibits stable and convergent training loss curves across both MedQA-USMLE and PubMedQA datasets, evidencing the robust training stability of the ranking-based preference optimization approach in medical reasoning tasks. Furthermore, the group-level ranking signals effectively mitigate the high variance issues inherent in pairwise preference methods.
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table 3 in Appendix A.2 provide representative cases generated by our model using the proposed prompt templates. These examples demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ 6 Conclusion and Future Work
In this paper, we present a novel approach for enhancing medical reasoning by combining ranked preference reinforcement optimization with probabilistic reasoning refinement. Our method introduces task-adaptive Chain-of-Thought generation, employs domain-specific reasoning templates, and incorporates a probabilistic quality assessment framework that enables targeted self-reflection and refinement. We evaluate our 1.1B parameter model on PubMedQA and MedQA-USMLE, demonstrating substantial improvements over much larger baseline models. These results validate that quality-driven reasoning enhancement is more effective than simple parameter scaling. Ablation studies further highlight the critical roles of reasoning refinement and ranked preference optimization.
In future work, we plan to extend our approach to additional medical reasoning tasks, integrate external medical knowledge bases, and explore optimal trade-offs between model scale and reasoning refinement quality. Moreover, conducting human evaluation studies with domain experts and real-world clinical assessments will provide valuable insights into practical utility and safety implications, ultimately advancing the deployment of AI in healthcare settings.
@@ A.1 Prompt Templates
Prompt Template (PubMedQA style)
You are a medical reasoning expert. Based on the question, context,
and final answer, produce a 4-step reasoning process:
1. Question Decomposition: Identify what the question is asking.
2. Background Knowledge: Recall relevant biomedical facts.
3. Logical Connection: Link knowledge to the question logically.
4. Final Justification: Conclude the answer with reasoning support.
Ensure factual consistency, no hallucination.
@@
Prompt Template (MedQA-USMLE style)
You are a clinical reasoning expert. Given the patient case and
final diagnosis, produce a 4-step diagnostic reasoning process:
1. Case Summary: Summarize the key patient information.
2. Clinical Significance: Explain the important findings.
3. Differential Diagnosis: Consider possible alternatives.
4. Most Likely Diagnosis: Justify the final diagnosis.
Ensure medical accuracy, no hallucination.

### v3

@@ 5.4 Evaluation Metrics
Main Results. Table 1 presents the performance comparison across all evaluated models under 0-shot, 1-shot, and 5-shot settings on both datasets. Our method consistently achieves strong gains across all configurations. On PubMedQA, RPRO (Gemma 2B) attains 60.23%, 61.76%, and 62.02% accuracy, with corresponding Macro F1 scores of 51.78%, 50.95%, and 53.12% in the 0-shot, 1-shot, and 5-shot settings, respectively. Compared to the best baseline (Gemma 7B), this represents improvements of up to 3.9% in accuracy and 3.8% in Macro F1, demonstrating the robustness of our refinement mechanism across different few-shot conditions. On MedQA-USMLE, our model achieves 42.38%, 46.27%, and 51.67% accuracy, with Macro F1 scores of 30.51%, 32.68%, and 35.24%. While maintaining competitive accuracy, RPRO consistently delivers higher Macro F1 than the strongest baseline (GPT-OSS 20B), with gains of +6.75%, +2.26%, and +3.13% across the three settings, highlighting better reasoning completeness and answer quality. Overall, our 2B-parameter model matches or surpasses much larger 7B–20B models, demonstrating the effectiveness and efficiency of our reasoning refinement strategy.
Clinical Evaluation on FEMH Dataset. To further verify the generalization and clinical applicability of our approach, we conducted an evaluation on the real-world Far Eastern Memorial Hospital (FEMH) clinical dataset. Each case consists of de-identified electronic medical records containing a clinician-written subjective description (S) and objective examination findings (O), along with the ground-truth diagnostic assessment (A). The model is required to generate the diagnostic statement based solely on the S+O inputs, and the generated assessment is compared against the physician-annotated ground truth.
We evaluate the semantic similarity between the model-generated diagnosis and the true assessment using BERTScore-F1 and sentence-level Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. As shown in Table 2, our proposed RPRO (Gemma 2B) achieves the best performance across all configurations, obtaining BERTScore-F1 / Cosine scores of 0.879 / 0.423, 0.884 / 0.507, and 0.891 / 0.528 in the respective settings. Compared with larger medical-tuned models such as BioMistral-7B and MedAlpaca-7B, RPRO demonstrates higher semantic coherence and diagnostic consistency, highlighting strong few-shot generalization and reliable clinical reasoning despite its compact model size.
Representative reasoning examples generated by our method on the FEMH dataset are shown in Table 4. These cases illustrate how the model integrates subjective symptoms and objective findings to produce structured diagnostic reasoning that aligns closely with clinician-provided assessments.
Ablation Studies. We conduct a comprehensive set of ablation experiments to assess how different components and key hyperparameters influence the behavior of RPRO. In addition to examining the effects of various training objectives and refinement strategies, we focus on three critical hyperparameters: the optimization weight $\beta$, which controls the balance between supervised alignment and preference-driven updates; the rollout budget $K$, which determines the diversity and stability of sampled reasoning trajectories; and the acceptance threshold, which governs the selectivity of the refinement process. By systematically varying these factors across the PubMedQA, MedQA-USMLE, and FEMH datasets, we evaluate their impact on convergence behavior, reasoning quality, and overall task performance.
Ablation on Training Objectives and Refinement. We begin our analysis by examining how different training objectives and refinement strategies contribute to overall performance. As shown in Table 3, standard supervised fine-tuning (SFT) with refinement provides a strong baseline, while DPO with refinement yields moderate improvements. RPRO without refinement already surpasses most baselines, demonstrating the advantage of groupwise ranking under the Bradley–Terry formulation. When refinement is added, RPRO achieves the best results across both PubMedQA and MedQA-USMLE, significantly improving both Accuracy and Macro F1. These results indicate that (1) the ranking-based preference optimization and (2) the reasoning refinement component are complementary and jointly responsible for the overall effectiveness of our method. This motivates the more fine-grained hyperparameter ablations presented in the following subsections.
Ablation on $\boldsymbol{\beta}$ Values. Fig. 2 presents the performance comparison under different $\beta$ values on the PubMedQA and MedQA-USMLE datasets. We vary $\beta\in\{0.0,0.05,0.1,0.2\}$ to control the balance between supervised alignment and preference-driven optimization. As shown in the figure, moderate $\beta$ values (around 0.1) yield the best results, achieving the highest Accuracy (60.23% on MedQA-USMLE, 42.38% on PubMedQA) and Macro F1 (55.83% and 30.51%, respectively). When $\beta$ is too small, the model underfits the preference signal; when $\beta$ is too large, over-regularization reduces reasoning diversity. This indicates that an appropriate trade-off between the base objective and preference regularization is crucial for stable optimization.
A similar trend is observed on the FEMH clinical dataset (right panel of Fig. 2), where $\beta=0.1$ achieves the best semantic alignment, with a BERTScore-F1 of 0.879 and a cosine similarity of 0.423. Both metrics degrade when $\beta$ deviates from this value, further confirming that moderate preference regularization leads to the most coherent and clinically meaningful diagnostic reasoning.
Ablation on Rollout Budget $\boldsymbol{K}$. To further investigate the influence of rollout budget on reasoning quality, we vary the number of rollouts $K\in{2,4,5,8}$ during preference optimization. As illustrated in Fig. 3, increasing $K$ from 2 to 5 leads to steady improvements in both Accuracy and Macro F1 on PubMedQA and MedQA-USMLE. However, when $K$ exceeds 5, performance gains saturate and even slightly decline, indicating that excessive rollouts introduce redundant samples and reward noise, increasing computational cost without meaningful benefit. Empirically, K=5K{=}5 provides the best trade-off between performance and efficiency, consistent with prior observations in RLHF and DPO frameworks where higher rollout counts yield diminishing returns.
On the FEMH clinical dataset (right panel of Fig. 3), a similar trend is observed. Semantic similarity metrics including BERTScore-F1 and cosine similarity increase as $K$ grows to 4 or 5, reaching peak values of 0.879 and 0.423 (shown as 0.88 and 0.42 in the figure due to rounding). When $K$ is further increased, both metrics slightly degrade, suggesting that excessive rollouts introduce noisy or redundant chains of thought and weaken alignment. These results confirm that a moderate rollout budget yields the most coherent and clinically meaningful diagnostic reasoning.
@@
Figure 2: Performance comparison on PubMedQA, MedQA-USMLE, and FEMH under different $\beta$ values. PubMedQA/MedQA report Accuracy (solid blue) and Macro F1 (dashed orange), while FEMH reports BERTScore-F1 (solid blue) and Cosine Similarity (dashed orange).
Figure 3: Performance comparison on PubMedQA, MedQA-USMLE, and FEMH under different rollout numbers ($K$). PubMedQA/MedQA report Accuracy and Macro F1; FEMH reports BERTScore-F1 and Cosine Similarity. Each prompt generates $K$ candidate CoTs. For $K<4$, all candidates are used ($M=K$); for $K\geq 4$, the top $M=4$ candidates are selected for training.
@@
Figure 4: Performance across different acceptance thresholds on PubMedQA, MedQA-USMLE, and FEMH. PubMedQA/MedQA report Accuracy (solid blue) and Macro F1 (dashed purple), while FEMH reports BERTScore-F1 and Cosine Similarity.
@@
Threshold and Convergency Analysis. Fig. 4 examines the impact of the acceptance probability threshold $\tau$ on model performance. Lower thresholds (0.2, 0.4) result in excessive refinement, potentially degrading reasoning quality through over-correction. The optimal threshold of 0.6 balances the necessity for refinement with the preservation of reasoning. Higher thresholds (0.8) show marginal accuracy improvements but decreased Macro F1, suggesting that conservative refinement may miss opportunities for quality enhancement while maintaining correctness.
@@
Figure 5: Training loss curves on MedQA-USMLE, PubMedQA, and FEMH. The plots show KL divergence, pairwise loss, ranking loss, and total loss across training steps for each dataset.
@@
To evaluate the robustness of our training procedure, we further analyze the loss dynamics of RPRO as shown in Fig. 5. The results demonstrate that RPRO exhibits stable and convergent loss curves on both MedQA-USMLE and PubMedQA, confirming that the ranking-based preference optimization framework maintains reliable optimization behavior throughout training. The group-level ranking objective also effectively mitigates the high-variance fluctuations typically observed in pairwise preference learning.
On the FEMH clinical dataset (right panel of Fig. 5), we observe a similarly stable convergence pattern. Both the pairwise and ranking components decrease smoothly over training steps, and the total loss reaches a consistent downward trajectory without oscillation. This indicates that RPRO also achieves stable optimization when applied to real-world clinical narratives, further validating its robustness and applicability beyond benchmark QA datasets.
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table 5 in Appendix A.2 provide representative cases generated by our model using the proposed prompt templates. In PubMedQA, the model explicitly decomposes the question, incorporates relevant biomedical background knowledge, and connects evidence to the final answer, reflecting a literature-grounded reasoning pattern. In MedQA-USMLE, the reasoning traces follow a clinical workflow, including case summarization, identification of key findings, differential diagnosis, and treatment selection, demonstrating the model’s ability to mimic expert diagnostic procedures. Across both tasks, the generated reasoning showcases coherent multi-step inference, consistent structure, and appropriate use of domain knowledge. Together, these examples show that our approach yields structured and interpretable reasoning that is well aligned with expert biomedical and clinical decision-making, offering qualitative evidence of the model’s improved reasoning quality.
@@ 6 Conclusion and Future Work
In this paper, we introduced a novel framework for enhancing medical reasoning by integrating ranked preference reinforcement optimization with probabilistic reasoning refinement. Our method incorporates task-adaptive Chain-of-Thought generation, domain-specific reasoning templates, and a probabilistic quality assessment mechanism that enables targeted self-reflection and iterative improvement. Through evaluations on PubMedQA, MedQA-USMLE, and the FEMH clinical reasoning dataset, our 2B-parameter model demonstrates performance gains that surpass those of substantially larger baseline models. These findings highlight that quality-driven reasoning enhancement can be more impactful than parameter scaling alone. Furthermore, ablation studies validate the importance of reasoning refinement and ranked preference optimization in achieving consistent and reliable improvements across diverse medical tasks.
Looking ahead, we aim to extend our approach to a broader set of medical reasoning benchmarks and incorporate structured external medical knowledge to further improve factual grounding. Another promising direction is exploring optimal trade-offs between model scale, reasoning quality, and computational efficiency. Additionally, conducting human evaluation studies with clinical experts and real-world deployment assessments will be essential for understanding practical utility, reliability, and safety considerations. Ultimately, we envision this work contributing toward the development of trustworthy, high-quality AI systems for clinical decision support.
7 Ethical approval
The study was approved by the Far Eastern Memorial Hospital (FEMH), Taiwan, Research Ethics Review Committee (FEMH IRB No.: 112086-F; https://www.femh-irb.org/ ), and all data were fully de-identified. All ethics review procedures and data collection were conducted in accordance with the committee’s standard guidelines and regulations (https://www.femh-irb.org/index.php/regulations ).
8 Funding
This work was supported by the Far Eastern Memorial Hospital Innovation Research Project.
9 Conflict of Interest
The authors declare that they have no competing interests or financial conflicts to disclose.
@@ A.1 Prompt Templates
Prompt Template (PubMedQA style) You are a medical reasoning expert. Based on the question, context, and final answer, produce a 4-step reasoning process: 1. Question Decomposition: Identify what the question is asking. 2. Background Knowledge: Recall relevant biomedical facts. 3. Logical Connection: Link knowledge to the question logically. 4. Final Justification: Conclude the answer with reasoning support. Ensure factual consistency, no hallucination.
@@
Prompt Template (MedQA-USMLE style) You are a clinical reasoning expert. Given the patient case and final diagnosis, produce a 4-step diagnostic reasoning process: 1. Case Summary: Summarize the key patient information. 2. Clinical Significance: Explain the important findings. 3. Differential Diagnosis: Consider possible alternatives. 4. Most Likely Diagnosis: Justify the final diagnosis. Ensure medical accuracy, no hallucination.
@@ B.1 PubMedQA
Figure 6: Examples of generated Chain-of-Thought (CoT) reasoning with varying quality levels on a PubMedQA sample. The high-quality CoT provides detailed, accurate, and non-redundant reasoning; the medium-quality example offers partial coverage; and the low-quality one demonstrates minimal reasoning coherence.
@@ B.2 MedQA-USMLE
Figure 7: Examples of generated Chain-of-Thought (CoT) reasoning at different quality levels. Each example shows the same question and gold answer but varies in reasoning coverage, factual accuracy, and redundancy.
@@ Appendix C Dataset Format (JSONL Example)
Figure 8: An illustrative example from the PubMedQA dataset showing question, context, and ranked reasoning breakdowns with multi-dimensional quality scores.